### v2

# Unified Sparse Mixture of Experts

Abstract
Sparse Mixture of Experts (SMoEs) models scale the capacity of models while maintaining constant computational overhead. Early designs typically relied on a fixed value of $k$, where $k$ represents either the number of experts selected per token or the number of tokens assigned per expert. However, these approaches encounter three key limitations: they may fail to route to important experts or tokens, may assign irrelevant ones, and often suffer from representation collapse among experts. This paper reexamines SMoEs through the lens of Linear Programming, and proposes a Unified Sparse Mixture of Experts (USMoE) framework that addresses these limitations. Specifically, our approach introduces a unified mechanism that integrates information from both the expert and token dimensions, and a unified scoring function that linearly combines similarity scores between experts and tokens. We provide both theoretical justification and empirical evidence demonstrating USMoE’s effectiveness in overcoming the limitations of traditional routing methods. Through comprehensive evaluations on both clean and corrupted settings for large language models and vision tasks, under both training-free and training scenarios, USMoE achieves up to a 10% performance improvement over standard approaches or reduces inference costs by up to 14%, while maintaining competitive accuracy.
1 Introduction
Sparse Mixture of Experts (SMoE) models have achieved notable success in natural language processing (NLP) and visual representation learning tasks (du_glam_2022, ; fedus_switch_2022, ; riquelme2021scalingvisionsparsemixture, ; shen-etal-2023-scaling, ). These advancements build on the Transformer architecture (NIPS2017_3f5ee243, ) and its variants (child2019generating, ; dai2019transformerxl, ), which leverage large datasets and significant compute resources. However, training large Transformer models can be prohibitively expensive, requiring extensive compute hours (kaddour2023challenges, ). To overcome this issue, SMoE models activate only a subset of experts for each input, reducing inference time compared to dense models (shazeer2017outrageously, ; zoph2022stmoe, ; artetxe2022efficient, ; krajewski2024scaling, ). The SMoE architecture can be categorized into two variants: Token Choice, which assigns experts to each token dai2024deepseekmoeultimateexpertspecialization ; qwen_moe ; muennighoff2024olmoeopenmixtureofexpertslanguage ; jiang2024mixtral , and Expert Choice, which assigns tokens to each expert zhou_mixture_experts_2022 . The advantage of Token Choice lies in its ability to dynamically select experts for each token, while Expert Choice ensures a more balanced token distribution across experts.
Input
USMoE
TC
EC
Input
USMoE
TC
EC
Figure 1: We compare token routing performance in vision tasks using 7×7 images, where each token is color-coded based on its assigned expert under setting $c=t$, where $c$ is a sparsity constraint and $t$ is number of image patches. Left: When the object is easy to distinguish, Expert Choice (EC) fails to assign different experts appropriately. Token Choice (TC) performs better but still does not align perfectly with the actual object, while USMoE correctly separates the object. Right: In more challenging images, both Expert Choice (EC) and Token Choice (TC) fail to distinguish between object and background. In contrast, USMoE successfully differentiates the object from the background, demonstrating greater efficiency in vision tasks compared to EC and TC, as further shown in Section 4.
Despite their promising results, SMoE models have several limitations. The Expert Choice approach suffers from token dropping  zhou_mixture_experts_2022 , while the Token Choice approach struggles with unbalanced expert loading  shazeer2017outrageouslylargeneuralnetworks . Additionally, both approaches are prone to representation collapse, where either a few experts dominate the routing or all experts learn similar representations chi_representation_2022 ; chen_towards_2022 . Recent research has explored improving router policies chi2022representation ; chen2023sparse ; do2023hyperrouter to mitigate these issues. However, existing methods face three key challenges: (1) The reliance on auxiliary losses requires careful balancing between the router loss and the task loss, which introduces trade-offs; (2) Token Choice (TC) struggles to handle noisy tokens effectively; and (3) Expert Choice (EC) suffers from information leakage issues, which significantly degrade performance on autoregressive models zhou2022mixtureofexpertsexpertchoicerouting ; wang2024auxiliarylossfreeloadbalancingstrategy ; raposo2024mixtureofdepthsdynamicallyallocatingcompute . As a result, the question of how to optimally select experts or tokens remains open.
Research Question What are the optimal criteria for selecting experts or tokens in SMoE models under fixed computational budgets?
Loading Balance Robustness Token Dropping Information Leak Expert Collapses
Token Choice Φ Φ Φ
Expert Choice Φ Φ
USMoE (ours) Φ Φ Φ Φ Φ
Table 1: Benefits of the proposed USMoEs’ model form over existing methods.
In this paper, we revisit SMoEs through the lens of Linear Programming NIPS2015_09662890 ; NEURIPS2021_a8fbbd3b to better understand the criteria for expert and token selection. From this perspective, Token Choice selects experts along the expert dimension by assigning each token to its most similar expert, while Expert Choice selects tokens along the token dimension by allowing each expert to choose the most similar tokens.
This perspective highlights a key trade-off: Expert Choice risks dropping important tokens, whereas Token Choice struggles to handle noisy or irrelevant tokens. Furthermore, both approaches suffer from the problem of representation collapse chi2022representation ; do2023hyperrouter ; pham2024competesmoe .
Building on this analysis, we propose the Unified Sparse Mixture of Experts (USMoE), a robust and efficient framework consisting of two key components: (1) the Unified Score and (2) the Unified Mechanism. The Unified Score is defined as a linear combination of two mapping functions that transform the dot product similarity between token and expert embeddings into a probability distribution - using functions such as softmax or sigmoid, as shown in Algorithm 1. In parallel, the Unified Mechanism incorporates information from both the token and expert dimensions, as illustrated in Figure 2. These components enable the SMoE model to dynamically prioritize tokens or experts while ensuring the selection of the most similar token-expert pair, enhancing both robustness and effectiveness. We summarize the benefits of USMoEs’ model form over existing methods in Table  1. To demonstrate the effectiveness of our approach, we evaluate USMoE across a variety of scenarios, including training-free, pretraining, and fine-tuning settings, spanning both language and vision domains under both clean and corrupted datasets. USMoE consistently outperforms baseline methods across these scenarios, with particularly strong gains in tasks that require deep input understanding, such as semantic textual similarity, classification, and clustering. Extensive experiments across various benchmarks show that USMoE achieves up to a 10% improvement over traditional approaches or reduces inference computational costs by 14%, all while maintaining high performance.
In summary, this paper makes the following key contributions:
  • We introduce a Linear Programming perspective on SMoE, highlighting the weaknesses of existing approaches.
  • We propose USMoE, a robust and efficient framework comprising two modules: (1) the Unified Score and (2) the Unified Mechanism, which together address the limitations of conventional approaches, as summarized in Table 1.
  • We theoretically demonstrate that USMoE effectively mitigates representation collapse, outperforming baseline methods.
  • We conduct extensive experiments on both large language models and vision domains, covering pretraining and both fine-tuning and training-free settings, providing a detailed analysis of USMoE’s performance and effectiveness.
2 Related Work
Sparse Mixture of Experts (SMoE). Sparse Mixture of Experts (SMoE), an extension of the Mixture of Experts framework (jacobs1991, ; jordan1994, ), has gained traction with large language models and has since been applied in various domains, including computer vision and speech recognition (NEURIPS2022_2f00ecd7, ; NEURIPS2021_48237d9f, ). The SMoE architecture consists of two main variants: Token Choice, where experts are assigned to each token (shazeer2017outrageouslylargeneuralnetworks, ; fedus_switch_2022, ; jiang2024mixtralexperts, ; do2024effectivenessdiscreterepresentationssparse, ), and Expert Choice, where tokens are assigned to specific experts (zhou_mixture_experts_2022, ).
Token Choice treats all tokens equally, which has raised concerns among researchers (wu2021attentionneed, ; hou2022tokendroppingefficientbert, ; lin2025rho1tokensneed, ), while Expert Choice suffers from token-dropping issues. Additionally, SMoE faces the challenge of representation collapse, where experts produce similar outputs. Various solutions have been proposed, such as XMoE, which employs low-dimensional routing scores (chi2022representation, ), and SMoE-dropout, which gradually activates more experts (chen2023sparse, ). Other approaches, including HyperRouter (do2023hyperrouter, ) and StableMoE (dai2022stablemoe, ), focus on enhancing router stability and robustness. Although these advancements have improved SMoE models, representation collapse remains a persistent issue (pham2024competesmoe, ; do2024simsmoesolvingrepresentationalcollapse, ). Our approach addresses this by optimizing the alignment between tokens and the most suitable experts, expanding expert specialization and mitigating collapse. Moreover, we also discuss related work on Linear Programming and other studies relevant to our approach in Section A.1.
3 Methodology
We introduce Unified Sparse Mixture of Experts (USMoE), a novel and efficient Sparse Mixture of Experts framework designed to address the limitations of both Token Choice and Expert Choice through a Unified Score, a scoring function that balances expert and token important and Unified Mechanism, a structured routing strategy that consider information from both token dimension and expert dimension.
3.1 Preliminaries
The Sparse Mixture of Experts (SMoE) architecture replaces the MLP layers in standard transformers with Mixture of Experts (MoE) layers (shazeer2017outrageously, ). Let $h\in\mathbb{R}^{T\times d}$ denote the token representations from the attention layer, where $T$ is the number of tokens and $d$ is the hidden dimension. Given $N$ expert functions $\{\text{FFN}_{j}\}_{j=1}^{N}$, a gating projection matrix $W\in\mathbb{R}^{d\times N}$ maps tokens to expert affinity scores. Denotes $S=f(hW)\in\mathbb{R}^{T\times N}$ is a compatibility score, where $f$ maps the scores to a routing distribution (e.g., softmax or sigmoid). Given $X\in\{0,1\}^{T\times N}$ be a binary routing matrix, where $x_{ij}=1$ if token $i$ is routed to expert $j$, the output of the SMoE layer is:
$f_{\text{SMoE}}(h)=\sum_{j=1}^{N}S_{\cdot j}\odot X_{\cdot j}\odot\text{FFN}_{j}(h\odot X_{\cdot j}),$ (1)
where $\text{FFN}_{j}(\cdot)$ denotes the computation performed by expert $j^{th}$, and $\odot$ represents element-wise multiplication, and FFN denotes for the Feed-forward neural networks.
3.2 Experts Selection as Linear Programming
Viewing expert selection as a linear programming problem allows for globally optimizing expert assignments by maximizing similarity scores under a strict sparsity (computational) constraint.
Expert Selection Optimization.  Given a computational budget of $c$ experts per token, the objective of Sparse Mixture-of-Experts (MoE) routing is to determine the optimal expert assignment $X^{*}$ that maximizes the total compatibility score, subject to a sparsity constraint $c$:
maximize $\sum_{i=1}^{T}\sum_{j=1}^{N}S_{ij}x_{ij}$ (2)
subject to $\sum_{i=1}^{T}\sum_{j=1}^{N}x_{ij}\leq c$
$x_{ij}\in\{0,1\},\quad\forall i,j.$
Figure 2: An illustration of our USMoE selection mechanism (middle) is shown, which incorporates information from both the token and expert dimensions. In contrast, Token Choice (TC, left) considers only the expert dimension, while Expert Choice (EC, right) focuses solely on the token dimension. Tokens marked with # represent noisy or irrelevant tokens. TC struggles to handle these # tokens effectively, and EC is prone to missing important tokens. USMoE addresses both issues, making our method more robust than traditional MoE approaches. This robustness is demonstrated in both theoretical analysis and experimental results. ’E’ denotes experts. Best viewed in color.
Proposition 3.1.
Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix and $c\in\mathbb{N}$ a global routing budget. Consider the objective:
$M=\langle S,X\rangle=\sum_{i=1}^{T}\sum_{j=1}^{N}S_{ij}x_{ij},$
subject to the constraint $\sum_{i,j}x_{ij}\leq c$, with $x_{ij}\in\{0,1\}$.
Let $X_{\text{USMoE}}=\text{TopK}(S,c)$ be the binary mask produced by selecting the top-$c$ entries of $S$ globally. Then for any other feasible binary routing matrix $X_{T}\in\{0,1\}^{T\times N}$ satisfying the same budget constraint $\sum_{i,j}(X_{T})_{ij}\leq c$, we have:
$\langle S,X_{T}\rangle\leq\langle S,X_{\text{USMoE}}\rangle.$
The Proposition 3.1 indicate that the Unified Mechanism (USMoE) yields the optimal solution that maximizes the total similarity under the global budget constraint. The proof is provided in Appendix  A.
Definition 3.2 (Unified Score Function).
Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix between $T$ tokens and $N$ experts. Define two score mapping functions:
  • $f_{t}(S)\in\mathbb{R}^{T\times N}$: a row-wise scoring function used in Token Choice (e.g., softmax applied across each row).
  • $f_{e}(S)\in\mathbb{R}^{T\times N}$: a column-wise scoring function used in Expert Choice (e.g., softmax applied across each column).
The Unified Score Function combines both perspectives via a linear combination:
$f_{\text{USMoE}}(S)=\alpha\cdot f_{e}(S)+\beta\cdot f_{t}(S),$
where $\alpha,\beta\in\mathbb{R}$ are non-negative coefficients such that $\alpha+\beta=1$.
The Unified Score as Definition 3.2 integrates both token-centric and expert-centric preferences to inform more balanced routing decisions.
Lemma 3.3.
The Unified Score Function $f_{\text{USMoE}}(S)=\alpha\cdot f_{e}(S)+\beta\cdot f_{t}(S)$, where $\alpha,\beta\in\mathbb{R}$ are non-negative coefficients such that $\alpha+\beta=1$, is more robust to representation collapse chi2022representation than using $f_{t}(S)$ or $f_{e}(S)$ alone.
Lemma 3.3 establishes that USMoE outperforms conventional SMoEs by mitigating the representation collapse issue, as discussed in chi2022representation . A detailed proof is presented in Appendix A.
3.3 Unified Sparse Mixture of Experts (USMoE)
Revising Expert Choice. The Expert Choice (EC) approach faces scalability challenges in autoregressive models due to information leakage wang2024auxiliarylossfreeloadbalancingstrategy ; raposo2024mixtureofdepthsdynamicallyallocatingcompute introduced by the softmax operator. To address this, we replace softmax with a sigmoid function for Token Choice, which eliminates inter-token leakage and, as shown in Figure 6, leads to improved bits-per-character (BPC) performance and a reduced token dropping ratio.
Finding 1: The Sigmoid function outperforms Softmax in Expert Choice MoE, as it handles information leakage and token dropping more effectively.
Unified Score. We empirically observe that Token Choice effectively captures in-context learning behavior, while Expert Choice conveys rich semantic alignment as Figure 3. To harness the complementary strengths of both, we propose the Unified Score, defined in Algorithm 1, as a weighted sum of the Token Choice and Expert Choice scores. This unified formulation integrates both token-centric and expert-centric signals to enable more balanced and informed routing decisions. Empirically, we find that setting the combination weight to $\alpha\approx 0.5$ yields a robust trade-off between the two strategies.
Figure 3: The performance of Unified Score (USMoE), Softmax (TC), and Sigmoid (EC) across the MTEB benchmark. Sigmoid outperforms Softmax on tasks without prompting, indicating stronger semantic representations.
Finding 2: Sigmoid (EC) excels at capturing semantic information, while Softmax (TC) is better suited for leveraging in-context cues.
Unified Mechanism. We propose the Unified Mechanism, which formulates expert-token assignment as a joint selection problem. By flattening the similarity matrix and selecting the top-$N$ expert-token pairs based on the highest unified scores, as illustrated in Figure 3.1, this mechanism enables efficient, context-aware routing within sparse MoE architectures.
[…]
2:Output $Y$
[…]
8:return $Y$
Algorithm 1 USMoE Layer
4 Experiments
In this section, we evaluate our method across both Large Language Models (LLMs) and Vision tasks, under clean and adversarial (attack) settings, and in both training-free and fine-tuned scenarios. We empirically demonstrate the advantages of USMoE over Token Choice (TC) and Expert Choice (EC) across advanced Sparse Mixture of Experts (SMoE) models, including QwenMoE qwen_moe , OLMoE muennighoff2024olmoeopenmixtureofexpertslanguage , and DeepSeekMoE dai2024deepseekmoeultimateexpertspecialization . Through extensive experiments, we show that: (1) USMoE outperforms baseline methods even without additional computational cost; (2) USMoE provides significant improvements across both language and vision domains; (3) Our method demonstrates robustness in both LLMs and vision tasks; and (4) USMoE supports flexible Top-k expert selection, which is valuable for scenarios with limited computational resources while maintaining competitive performance.
4.1 Large Language Models (LLMs)
We evaluate our method against two standard SMoE baselines across the following three settings:
Training-Free.  We evaluate our approach on three state-of-the-art Sparse Mixture of Experts (MoE) models: (1) OLMoE-1B-7B muennighoff2024olmoeopenmixtureofexpertslanguage , which has 7B parameters, 16 layers, and 64 experts per layer; (2) Qwen1.5-MoE-A2.7B qwen_moe , comprising 7B parameters, 24 layers, and 60 experts per layer; and DeepSeekMoE-16B which consists of 16B parameters, 28 layers, and 64 experts per layer. We evaluate these models on the Massive Text Embedding Benchmark (MTEB) muennighoff2023mtebmassivetextembedding without additional fine-tuning.
Supervised Fine-Tuning.  We demonstrate the effectiveness of USMoE on supervised fine-tuning using the Alpaca dataset alpaca , adopting the memory-efficient full-parameter training approach proposed by zhao2024galorememoryefficientllmtraining , which outperforms common low-rank adaptation methods in terms of memory efficiency. To evaluate the robustness of our method, we follow the procedure in nielsen2025tightclustersmakespecialized , applying a Text Attack strategy where words are randomly replaced with a generic token “AAA”.
Training from scratch.  To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE jiang2024mixtralexperts , SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE chi2022representation , and StableMoE dai2022stablemoe , as well as the Expert Choice approach zhou2022mixtureofexpertsexpertchoicerouting for pre-training and fine-tuning tasks. We follow the approach of (chen_sparse_2023, ) and use a base Transformer-XL dai_transformer_xl_2019 with four decoder layers. We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in chen_sparse_2023 . Then we fine-tune the pre-trained weights for text classification tasks, including SST-2 socher_recursive_2013 , SST-5 socher_recursive_2013 , IMDB maas_learning_2011 , and BANKING77 casanueva-etal-2020-efficient . More implementation details and additional results are provided in the Appendix A.
4.1.1 Training-Free
In this section, inspired by  li2024mixtureofexpertsllmsecretlyembedding , we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B muennighoff2024olmoeopenmixtureofexpertslanguage :, DeepSeekMoE-16B  dai2024deepseekmoeultimateexpertspecialization , Qwen1.5-MoE-A2.7B qwen_moe . We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  muennighoff2023mtebmassivetextembedding , which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization. Following the MTEB evaluation framework, we use Accuracy for Classification, V-Measure for Clustering, Average Precision for Pair Classification, Mean Average Precision for Re-ranking, nDCG for Retrieval, and Spearman’s correlation for STS and Summarization. Our method consistently demonstrates performance improvements across a range of MTEB tasks . Detailed results for datasets under each task type are provided in Appendix A. USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
For tasks evaluated in Figure  4, USMoE proves even more effective at enhancing the Token Choice approach, delivering notable gains of 14%, 12%, and 13% for OLMoE-1B-7B, Qwen1.5-MoE-A2.7B, and DeepSeekMoE-16B, respectively, across MTEB tasks without additional training. Specifically, Qwen1.5-MoE-A2.7B achieves a remarkable improvement from 13.4% (Token Choice) to 40.0% (USMoE) in the Summarization task, representing a 198% gain. This trend persists across DeepSeekMoE-16B and OLMoE-1B-7B, where USMoE consistently outperforms both the Token Choice and Expert Choice approaches. Overall, our approach outperforms the baselines in terms of performance while exhibiting lower variance across multiple tasks and different runs.
Figure 4: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks and advance SMoE models. The best result for each row is highlighted in bold. Best viewed in color.
4.1.2 Supervised Fine-Tuning
Supervised Fine-Tuning.  We demonstrate the efficiency and robustness of USMoE in supervised fine-tuning across two advanced SMoE models: QwenMoE (7B)qwen_moe and OLMoE (7B)muennighoff2024olmoeopenmixtureofexpertslanguage . For training, we adopt Supervised Fine-Tuning (SFT) using GaLorezhao2024galorememoryefficientllmtraining , which enables full-parameter optimization with improved memory efficiency compared to typical low-rank adaptation methods such as LoRA. The fine-tuning is conducted on the Alpaca datasetalpaca for 2,000 steps, with evaluation performed every 10 steps. Furthermore, we assess our method on both clean (referred to as QwenMoE and OLMoE) and adversarially corrupted versions (denoted QwenMoE-Corrupt and OLMoE-Corrupt), where sequences of random “AAA” tokens are injected to simulate noise.
Robustness.  Figure 5 presents the training results for both QwenMoE and OLMoE models. The results show that USMoE consistently outperforms the baselines across both QwenMoE and OLMoE models. Our method surpasses both Token Choice (TC) and Expert Choice (EC), particularly on QwenMoE, where it achieves significantly lower training and validation perplexity, indicating faster convergence and better generalization. Notably, the performance gap becomes even more pronounced under the corrupted setting, highlighting the superior robustness of USMoE compared to traditional SMoE approaches.
Figure 5: Illustration of comparing the performance of USMoE, Token Choice (TC), Expert Choice (EC) using QwenMoE and OLMoE for the Supervised Fine-Tuning task on Apaca dataset for 2K steps under both clean and corrupted settings. Training and validation perplexity over training steps are reported, and lower values are better.
4.1.3 Training from Scratch
Robustness.  Table 2 presents pre-training results on four datasets (enwik8, text8, WikiText-103, and One Billion Words) under both clean and corrupted settings. USMoE consistently outperforms both Expert Choice and Token Choice routing methods across all datasets. Similar to the SFT task, the performance gap widens under corruption, especially on One Billion Words—highlighting the robustness advantage of USMoE over traditional SMoEs. For additional experiments on training from scratch, such as scaling models up to 400M parameters or fine-tuning on downstream tasks—please refer to Section A.
Flexibility.  One of USMoE’s key strengths is its flexibility in computational resource usage. Unlike traditional SMoEs, which require the number of selected experts (top-$k$) to be an integer, USMoE supports fractional expert selection. This allows the model to maintain competitive performance while leveraging fewer experts. For instance, on text8, USMoE outperforms SMoE using only 1.5 experts. It also surpasses Token Choice, which uses two experts, on enwik8, text8, and One Billion Words, while averaging just 1.5 experts per token. Moreover, using only 1.5 experts enables USMoE to reduce FLOPs by 14% compared to SMoE, with minimal performance trade-off. This makes USMoE particularly well-suited for resource-constrained environments such as IoT and edge devices.
Setting Original Corrupt
Transformer-XL(20M) Topkk Enwik8 Text8 WikiText-103 lm1b Enwik8 Text8 WikiText-103 lm1b
USMoE 2 1.18 1.20 29.20 56.90 1.75 1.83 38.45 68.43
1.5 1.19 1.28 30.67 57.55 1.78 1.95 40.39 70.47
TC 2 1.20 1.29 30.16 58.00 1.77 1.86 39.31 79.73
EC 2 1.18 1.24 29.83 58.60 1.76 1.89 39.28 71.75
Table 2: Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with BPC on the Enwik8 and Text8 test sets, and perplexity on the WikiText-103 and One Billion Word test sets. Lower values are better, with the best results highlighted in bold.
4.2 Vision
Effiency. To ensure a thorough and meaningful performance comparison, we evaluate against two categories of baselines that are closely aligned with USMoE, including Token Choice, Expert Choice, and SoftMoE (puigcerver2024sparsesoftmixturesexperts, ), currently among the most advanced MoE models in the vision domain. We explore two Mixture of Experts configurations: (1) a small model with 10 million parameters (10M), and (2) a large model with 110 million parameters (110M). As shown in Table 3, USMoE consistently outperforms Vision Transformer variants with Token Choice, Expert Choice, and SoftMoE across all eight tasks and four image classification datasets. Our experiments are conducted three times on four datasets (CIFAR-10, CIFAR-100, STL-10, SVHN, and ImageNet1K), using different random seeds. We report average results along with standard deviations. USMoE not only achieves superior average performance but also demonstrates greater stability, as evidenced by lower standard deviation values.
ViT-MoE 10M 110M Avg.
Dataset Cifar10 Cifar100 STL-10 SVHN ImageNet-1K Cifar10 Cifar100 STL-10 SVHN ImageNet-1K
USMoE 89.6±0.3 66.6±0.5 66.7±0.4 95.6±0.1 60.2±0.1 91.5±0.5 67.3±0.5 66.2±0.1 96.1±0.1 73.5±0.4 77.3±0.3
TC 88.7±0.2 65.4±0.5 66.4±0.1 95.1±0.1 56.6±0.5 85.7±0.8 55.5±2.8 64.4±0.2 94.5±0.4 72.0±0.4 74.4±1.5
EC 88.9±0.3 65.7±0.4 66.1±0.4 95.0±0.1 56.2±0.4 87.4±0.7 66.2±0.9 65.5±0.4 93.2±0.2 70.9±0.5 75.5±1.2
SoftMoE 85.6±0.3 61.4±0.3 65.4±0.2 94.8±0.1 46.8±0.6 80.3±1.0 42.9±1.4 63.9±1.2 93.5±0.1 71.2±0.3 70.6±1.5
Table 3: Accuracy of models evaluated on vision datasets. Each method is evaluated 3 times, reporting the mean and standard deviation. Higher is better, the best results are in bold.
Robustness.  Table 4 presents a comparative analysis of ViT-MoE variants, USMoE, Token Choice (TC), Expert Choice (EC), and SoftMoE, under three common adversarial attacks: Projected Gradient Descent(PGD)(chang2018efficienttwostepadversarialdefense, ), Fast Gradient Sign Method(FGSM) goodfellow2015explainingharnessingadversarialexamples , and Simultaneous Perturbation Stochastic Approximation (SPSA) (SPALL1997109, ), across five standard image classification datasets. USMoE consistently achieves strong performance, outperforming or matching the best baselines in most scenarios. Notably, it achieves the highest average robustness across all three attack types, particularly excelling under PGD (46.2%), FGSM (43.3%), and SPSA (63.9%). While Token Choice (TC) performs poorly across all three attack types, USMoE consistently delivers stronger and more reliable performance on average. These results highlight USMoE’s robustness and generalization capability under adversarial conditions, making it a more stable and effective choice for secure image classification tasks.
Finding 3: Token Choice (TC) struggles under adversarial attacks.
Dataset PGD FGSM SPSA
ViT-MoE (10M) USMoE TC EC SoftMoE USMoE TC EC SoftMoE USMoE TC EC SoftMoE
CIFAR‑10 57.4 27.7 57.5 55.6 49.5 26.4 48.5 45.8 83.6 33.4 87.0 69.1
CIFAR‑100 27.5 11.9 27.0 28.6 21.5 11.1 19.3 20.6 65.8 15.8 59.1 39.3
STL‑10 41.0 37.7 39.5 39.8 36.5 35.8 34.8 35.0 64.3 44.1 61.3 49.6
SVHN 81.0 38.4 80.5 91.3 76.0 36.1 75.4 69.2 92.5 36.3 91.8 84.4
ImageNet‑1K 24.0 1.3 22.4 13.3 12.2 1.4 11.2 9.9 13.4 1.5 12.3 10.9
Avg. 46.2 23.4 45.4 45.7 43.3 22.2 37.8 36.1 63.9 26.2 62.3 50.7
Table 4: Robustness evaluation of different ViT-MoE models under adversarial attacks: PGD, FGSM, and SPSA across five datasets. Bold indicates the best performance for each task and attack.
5 Conclusion
In this work, we reinterpret Token Choice and Expert Choice Sparse Mixture of Experts (SMoE) through a linear programming lens, revealing their inherent limitations. Based on this analysis, we propose Unified Sparse Mixture of Experts (USMoE) - a novel and efficient framework that advances SMoE by introducing a unified mechanism and unified score learning. We theoretically prove that USMoE achieves superior expert selection compared to traditional methods, effectively improving expert learning capacity while mitigating expert collapse. As a result, USMoE learns more robust expert representations and overcomes the representation collapse issues commonly observed in conventional SMoE training. Extensive experiments across diverse domains, including vision and large language models (LLMs), under both clean and adversarial settings (covering training-free and fine-tuning scenarios), demonstrate that USMoE enables more efficient and effective training and inference compared to state-of-the-art routing strategies.
Appendix A Appendix
Supplementary Material for “Unified Sparse Mixture of Experts”
This document is organized as follows: Appendix A.1 discuss related work on Linear Programming and other studies relevant to our approach. We also discuss more about the Ablation Studies at Appendix A.2. Appendix A.3 provides the theoretical proof supporting the discussion in Section 3. Following this, Appendix A.4 additional experiment results that show the advantage of our method. Appendix A.5 offers a detailed analysis of the Token Choice and Expert Choice routers and explains why our method outperforms the baseline approaches. Appendix A.6 describes the implementation details in full.
A.1 Related Work
Linear Programming. Linear programming, a foundational method in mathematical optimization, has gained relevance in the domain of Large Language Models (LLMs), where task allocation and dependency management among skills can be effectively modeled as constrained optimization problems [40, 33]. Recent work such as [22] introduces an Integer Linear Programming (ILP) formulation to jointly optimize expert placement by accounting for token load, communication overhead, and computation cost. Framing SMoEs from the perspective of linear programming not only offers deeper theoretical understanding but also opens promising directions for developing more efficient and principled routing strategies.
The most relevant work to ours is Expert Race[56], which proposes a flexible routing strategy for diffusion transformers. However, similar to the traditional Expert Choice (EC) approach, it suffers from a critical limitation, information leakage - where a model unintentionally accesses future tokens [43, 54], which becomes especially problematic in autoregressive language modeling settings[43, 54]. In contrast, we propose a modified version of Expert Choice [58] by replacing the softmax mapping function with a sigmoid function and limiting comparisons to within the token dimension. This design effectively mitigates information leakage, making Expert Choice more suitable for large language models.
Additionally, we introduce two novel modules: (1) the Unified Score, a linear combination of the mapping functions from Token Choice and Expert Choice, and (2) the Unified Mechanism, which jointly considers both expert and token dimensions when selecting experts. Both theoretical analysis and empirical results demonstrate that our method outperforms traditional SMoE approaches in terms of both performance and robustness.
A.2 Ablation Studies
We investigate the effectiveness and robustness of USMoE to the different hyper-parameter settings.
A.2.1 Unified Learning Strategy Comparison
The Unified Mechanism is implemented using two approaches: (1) a sequence-based method that compares all tokens within a sequence (referred to as "USMoE-Sequence") and (2) a batch-based method that compares all tokens within a batch or mini-batch (referred to as "USMoE-Batch"). We evaluate both approaches on the Classification task, with results presented in Table 10. The findings indicate that both methods outperform the Expert Choice and Token Choice approaches, demonstrating the effectiveness of our method. Notably, the sequence-based approach achieves superior performance in the Classification task, as it ensures that no important tokens are missed within a sequence - an assurance that the batch/mini-batch implementation may not always provide.
A.2.2 Robustness to the controlling factor
The Unified Score ($\alpha$) enables the model to adjust its scoring mechanism, either favoring a diverse set of experts per sequence or distributing experts more evenly across tokens. We evaluate the robustness of the controlling factor $\alpha$ on the classification task using the DeepSeekMoE-16B model, with results presented in Table 11. When $\alpha=0.0$, the scoring mechanism aligns with Token Choice, while at $\alpha=1.0$, it follows Expert Choice.
Overall, USMoE demonstrates strong effectiveness within the range of $\alpha\in(0.3,0.7)$, striking a balance between expert diversity and token importance. This range provides an optimal trade-off between enforcing SMoE’s unified policy and enhancing traditional approaches for the task, leading to superior overall performance. Notably, all tested $\alpha$ configurations outperform the Expert Choice approach ($\alpha=1.0$).
A.3 Theoretical Proof for Section  3
Definition A.1 (General Routing Function $\mathcal{R}$).
Let $S\in\mathbb{R}^{T\times N}$ be the compatibility score matrix between $T$ tokens and $N$ experts, and let $c\in\mathbb{N}$ denote a global routing budget that constrains the total number of token-to-expert assignments. Let $\text{TopK}_{R}$, $\text{TopK}_{C}$, and TopK denote the TopK selection operators applied row-wise, column-wise, and globally, respectively. We define a general routing function $\mathcal{R}(S,c,\texttt{mode})$ that constructs a sparse binary routing matrix $X\in\{0,1\}^{T\times N}$, where $x_{ij}=1$ indicates that token $i$ is routed to expert $j$, as follows:
$X=\mathcal{R}(S,c,\texttt{mode})=\begin{cases}\text{TopK}(S,c),&\texttt{mode}=\text{USMoE},\\[6.0pt] \text{TopK}_{R}(S,\left\lfloor\frac{c}{T}\right\rfloor),&\texttt{mode}=\text{Token Choice},\\[6.0pt] \text{TopK}_{C}(S,\left\lfloor\frac{c}{N}\right\rfloor),&\texttt{mode}=\text{Expert Choice},\end{cases}$
Lemma A.2.
The Unified Mechanism (USMoE), as defined in Definition A.1, yields a superior solution to the optimization problem in Equation 2 compared to both Token Choice and Expert Choice mechanisms.
The proof is given in Appendix A, where Lemma A.2 establishes the superiority of USMoE over Token Choice and Expert Choice.
A.3.1 Proof of Lemma  A.2
Let the ordered arrays $(S^{t}_{1},S^{t}_{2},\dots,S^{t}_{T})$, $(S^{e}_{1},S^{e}_{2},\dots,S^{e}_{T})$, and $(S^{u}_{1},S^{u}_{2},\dots,S^{u}_{T})$ represent the top-$T$ compatibility scores selected by the Token Choice, Expert Choice, and Unified Score mechanisms, respectively. These scores are sorted in ascending order such that
$S^{t}_{1}\leq S^{t}_{2}\leq\dots\leq S^{t}_{T},\quad S^{e}_{1}\leq S^{e}_{2}\leq\dots\leq S^{e}_{T},\quad S^{u}_{1}\leq S^{u}_{2}\leq\dots\leq S^{u}_{T}.$
We claim the following inequality holds:
$S^{t}_{i}$ $\leq S^{u}_{i},\quad\forall i\in[1,T],$ (3)
$S^{e}_{i}$ $\leq S^{u}_{i},\quad\forall i\in[1,T].$
Proof.
Let $x\in\mathbb{R}^{T\times d}$ denote a matrix of token embeddings and $\mathcal{E}\in\mathbb{R}^{d\times N}$ the matrix of expert embeddings, where $d$ is the embedding dimension, $T$ is the number of tokens, and $N$ is the number of experts. The similarity matrix $S\in\mathbb{R}^{T\times N}$ between tokens and experts is computed via a dot product:
$S=x\cdot\mathcal{E}.$ (4)
Since the softmax function is monotonic, the top-$k$ selection remains invariant under softmax transformation:
$\operatorname{TopK}(\operatorname{softmax}(S),k)=\operatorname{TopK}(S,k).$ (5)
For the Token Choice approach, the top similarity score for each token is defined as:
$S^{t}_{j}=\max_{k\in[1,N]}S_{jk},$ (6)
and across all tokens, the top-$T$ scores are aggregated into $\{S^{t}_{i}\}_{i=1}^{T}$.
In contrast, the Unified Mechanism selects the top-$T$ scores globally across all token-expert pairs:
$S^{u}_{j}=\max_{(q,k)\in[1,T]\times[1,N]}S_{qk}.$ (7)
By construction, it follows that:
$S^{t}_{i}\leq S^{u}_{i},\quad\forall i\in[1,T],$ (8)
since Token Choice selects top experts per token independently, while Unified Score considers all token-expert pairs jointly.
Similarly, in the Expert Choice approach, each expert selects its most compatible token:
$S^{e}_{i}=\max_{q\in[1,T]}S_{qi},\quad\forall i\in[1,N],$ (9)
and again, the top-$T$ scores across experts are extracted as $\{S^{e}_{i}\}_{i=1}^{T}$. As the Unified approach considers the global maximums, we have:
$S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,T].$ (10)
Combining (8) and (10), we conclude:
$S^{t}_{i}\leq S^{u}_{i},\quad S^{e}_{i}\leq S^{u}_{i},\quad\forall i\in[1,T],$ (11)
which completes the proof. ∎∎
A.3.2 Proof of Proposition 3.1
Proof.
Following  [9] and  [18], we illustrate the representation collapse issue using the Jacobian matrix approach. Specifically, the Jacobian matrix of the SMoE with respect to $x\in\mathbb{R}^{b\times d}$ is given as:
$\boldsymbol{J_{SMoE}}=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}+\sum_{j=1}^{n}\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\boldsymbol{E}(x)_{i}\boldsymbol{e}_{j}^{\top}$
$\Longrightarrow\boldsymbol{J_{SMoE}}=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}+\sum_{j=1}^{n}\boldsymbol{c}_{j}\boldsymbol{e}_{j}^{\top},$ (12)
where $\boldsymbol{c}_{j}=\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\boldsymbol{E}(x)_{i}$. The first part of Equation  12, $\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}$, represents the contribution from the input token and experts to the final output. The second part, (2) $\sum_{j=1}^{n}\boldsymbol{c}_{j}\boldsymbol{e}_{j}^{\top}$ relates to learning an improved gating function to minimize task loss. Furthermore, Equation  12 should be updated as a linear combination of expert embeddings. Due to $n<<d$ in practice, the above equation illustrates the representation.
Given $\mathcal{S}(x)=\alpha\times\mathcal{S}_{e}(x)+(1-\alpha)\times\mathcal{S}_{t}(x)$, where $\mathcal{S}_{e}(x)$ and $\mathcal{S}_{t}(x)$ represent the similarity score functions for expert choice and token choice, respectively, the Jacobian matrix of USMoE with respect to $x\in\mathbb{R}^{b\times d}$ is expressed as:
$\boldsymbol{J_{U}}$ $=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}}+\sum_{j=1}^{n}\mathcal{S}_{e}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{e}(x)_{j}\right)\boldsymbol{E}(x)_{k}\boldsymbol{e}_{j}^{\top}$
$\quad+\sum_{j=1}^{n}\mathcal{S}_{t}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{t}(x)_{j}\right)\boldsymbol{E}(x)_{k}\boldsymbol{e}_{j}^{\top}.$
$\Longrightarrow\boldsymbol{J_{U}}=J_{1}+\sum_{j=1}^{n}c_{j}e_{j}^{\top}+\sum_{j=1}^{n}d_{j}e_{j}^{\top}$ (13)
where
$J_{1}=\mathcal{S}(x)_{k}\boldsymbol{J}^{\mathrm{FFN}},$ (14)
$c_{j}=\mathcal{S}_{e}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{e}(x)_{j}\right)\boldsymbol{E}(x)_{k}\boldsymbol{e}_{j}^{\top},$ (15)
$d_{j}=\mathcal{S}_{t}(x)_{k}\left(\delta_{kj}-\mathcal{S}_{t}(x)_{j}\right)\boldsymbol{E}(x)_{k}.$ (16)
Similar to the Jacobian matrix of SMoE as Equation  12, the Jacobian matrix of USMoE also consists of two terms: (1) $J_{1}$, which depends on the input token and experts for the final output; and (2) $\sum_{j=1}^{2n}o_{j}\boldsymbol{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss. Since $2n>>n$, USMoE is more effective than SMoE in addressing the representation collapse issue.
A.3.3 Proof of Lemma 3.3
Proof.
Let $S\in\mathbb{R}^{t\times n}$ be the similarity matrix, and let $c\in\mathbb{N}$ be a global routing budget. The goal is to find a binary matrix $X\in\{0,1\}^{t\times n}$ such that
$\sum_{i=1}^{t}\sum_{j=1}^{n}x_{ij}\leq c$
and the total score
$M=\langle S,X\rangle=\sum_{i=1}^{t}\sum_{j=1}^{n}S_{ij}x_{ij}$
is maximized.
Let $X_{\text{USMoE}}=\text{TopK}(S,c)$ denote the binary matrix that marks the positions of the top-$c$ largest entries in $S$. By construction, $X_{\text{USMoE}}$ sets exactly $c$ entries of $S$ with the highest values to 1 and all others to 0.
Suppose there exists another feasible solution $X_{T}\in\{0,1\}^{t\times n}$ with
$\sum_{i,j}(X_{T})_{ij}\leq c$
and assume for contradiction that
$\langle S,X_{T}\rangle>\langle S,X_{\text{USMoE}}\rangle.$
Since $X_{T}$ selects at most $c$ entries from $S$, and the values of the selected entries are summed to compute $\langle S,X_{T}\rangle$, the only way for $\langle S,X_{T}\rangle$ to exceed $\langle S,X_{\text{USMoE}}\rangle$ is if some entries selected by $X_{T}$ are larger than those selected by $X_{\text{USMoE}}$. However, this contradicts the definition of $X_{\text{USMoE}}$ as containing the top-$c$ largest values in $S$.
Therefore, for any feasible $X_{T}$, we must have
$\langle S,X_{T}\rangle\leq\langle S,X_{\text{USMoE}}\rangle,$
which proves that $X_{\text{USMoE}}$ is the optimal solution. ∎
A.4 Experimental details and additional results
A.4.1 Which Mapping Function for Expert Choice?
The Expert Choice (EC) approach faces scalability challenges for large language models due to the issues of information leakage [54, 43] and token dropping [42]. We revisit this problem and identify that these issues stem from the use of the softmax operator, which introduces information leakage across tokens, and the top-k selection applied along the batch dimension.
To address this, we propose replacing the softmax mapping function with a sigmoid function to eliminate inter-token information leakage. Additionally, we mitigate batch-wise leakage by applying the top-k selection only along the sequence dimension.
Figure 6 compares the performance of EC using softmax versus sigmoid. The results show that the sigmoid-based EC not only achieves better bits-per-character (BPC) performance but also reduces the token dropping ratio. This improvement suggests that using the sigmoid mapping function has strong potential for scaling up the EC approach in large language models.
Figure 6: Performance comparison of the Expert Choice approach using two mapping functions: sigmoid and softmax, evaluated on training performance and token dropping ratio (lower is better).
A.4.2 Training-free additional results
In this section, we compare methods both with and without prompts, including PromptEOL [29]. Inspired by  [34], we also compare our method with an approach that uses the similarity score between router and expert embeddings as the hidden representation, which we refer to as "Router Embedding" or simply "Router" Additionally, we evaluate against MoEE  [34], which leverages both Router Embedding and the hidden representation of the SMoE model as embeddings.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 43.1 57.7 56.2 51.7 61.4
Clustering 16.2 24.8 26.9 23.2 32.1
PairClassification 53.5 62.0 58.9 66.0 68.9
Reranking 41.7 51.3 51.0 53.2 55.1
STS 49.4 63.5 44.2 67.8 71.1
Summarization 25.6 28.9 29.7 30.4 30.5
Average 38.3 48.0 44.5 48.7 53.2
Qwen1.5-MoE-A2.7B Classification 48.8 58.0 35.2 54.0 59.7
Clustering 14.3 34.2 29.2 30.1 37.5
PairClassification 51.9 60.5 56.0 60.3 66.6
Reranking 41.0 46.6 45.0 51.1 56.8
STS 48.3 50.1 50.0 64.3 69.0
Summarization 27.0 23.0 21.9 27.3 31.0
Average 38.6 45.4 39.6 47.9 53.4
DeepSeekMoE-16B Classification 48.6 56.4 55.4 53.0 60.4
Clustering 17.8 29.0 20.3 28.5 32.8
PairClassification 57.4 59.8 53.8 63.3 67.9
Reranking 43.8 45.7 40.9 50.6 52.4
STS 52.8 49.0 37.1 63.4 68.1
Summarization 29.1 24.4 25.7 29.2 30.7
Average 41.6 44.0 38.9 48.0 52.0
Table 5: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL [28]. The best result for each row is highlighted in bold.
Model Task Router TC EC MoEE USMoE
OLMoE-1B-7B Classification 41.2 43.4 44.9 41.8 51.3
Clustering 13.7 14.7 12.0 14.5 21.0
PairClassification 45.3 39.1 35.5 45.7 61.7
Reranking 37.5 37.4 35.3 39.5 46.5
STS 39.9 24.1 18.2 39.9 59.6
Summarization 28.4 20.9 21.1 29.8 31.8
Average 34.3 29.9 27.8 35.2 45.3
Qwen1.5-MoE-A2.7B Classification 43.8 50.3 25.5 47.7 52.2
Clustering 13.6 27.4 23.2 25.2 29.9
PairClassification 45.9 46.9 43.4 51.5 56.2
Reranking 39.6 45.3 41.6 48.5 53.1
STS 38.8 38.0 35.6 51.8 60.7
Summarization 28.3 13.4 15.1 31.2 40.0
Average 35.0 36.9 30.7 42.6 48.7
DeepSeekMoE-16B Classification 43.4 46.6 44.7 44.4 49.4
Clustering 13.4 18.1 13.5 17.8 21.0
PairClassification 45.5 40.9 37.1 46.1 53.5
Reranking 38.5 38.9 35.1 42.2 45.7
STS 37.7 26.3 23.3 40.2 51.2
Summarization 24.9 22.0 18.5 24.4 29.9
Average 33.9 32.1 28.7 35.9 41.8
Table 6: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks without prompts and models. The best result for each row is highlighted in bold.
In this section, inspired by  [34], we test our method as a plug-in framework on well-trained SMoE models, including OLMoE-1B-7B [37]:, DeepSeekMoE-16B  [12], Qwen1.5-MoE-A2.7B [52]. We evaluate performance on a subset of tasks from the Massive Text Embedding Benchmark (MTEB)  [38], which covers key downstream applications for sentence embeddings, including Classification, Clustering, Pair Classification, Re-ranking, Retrieval, Semantic Textual Similarity (STS), and Summarization. Following the MTEB evaluation framework, we use Accuracy for Classification, V-Measure for Clustering, Average Precision for Pair Classification, Mean Average Precision for Re-ranking, nDCG for Retrieval, and Spearman’s correlation for STS and Summarization. Our method consistently demonstrates performance improvements across a range of MTEB tasks in two scenarios: (1) with PromptEOL [28] (denoted as "with prompts" for brevity), as shown in Table 5, and (2) without prompts, as shown in Table 6. Detailed results for datasets under each task type are provided in Appendix A. USMoE outperforms both the Expert Choice and Token Choice approaches in most cases, underscoring the complementary nature of these two methods.
Interestingly, Router Embedding is less affected by prompting on the Classification dataset, as shown in Figure 8(a), while Token Choice, Expert Choice, and USMoE (ours) achieve significant performance improvements in the prompting setting. Figure 8(a) also demonstrates that our method is not only more effective but also more stable than the baselines, as indicated by a lower standard deviation. Additionally, Figure 8(b) illustrates the distribution of our method across MTEB tasks in both prompted and non-prompted scenarios. Overall, our approach outperforms the baselines in terms of performance while exhibiting lower variance across multiple tasks and different runs.
We provide a detailed evaluation of three state-of-the-art SMoE models: OLMoE-1B-7B (Table  7), Qwen1.5-MoE-A2.7B (Table 8), and DeepSeekMoE-16B (Table  9). Our results demonstrate the effectiveness of our method across various models and prompts, comparing its performance against baseline approaches such as Token Choice (TC) and Expert Choice (EC).
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification OLMoE Emotion None 24.1 24.5 26.3 25.1 35.8
Prompt 27.6 49.9 49.0 44.5 54.8
Toxic None 51.9 58.9 59.9 51.9 62.8
Prompt 52.3 65.2 61.3 53.4 67.6
Tweet None 47.7 46.8 48.3 48.4 55.2
Prompt 49.5 58.0 58.4 57.2 61.7
Clustering OLMoE Medrxiv None 15.0 17.6 14.8 17.4 20.6
Prompt 15.8 23.9 27.7 22.0 28.2
20Groups None 12.4 11.8 9.2 11.5 21.3
Prompt 16.7 25.7 26.2 24.4 36.0
Pair Classification OLMoE SemEval None 43.6 35.8 31.3 43.6 50.2
Prompt 45.7 46.7 40.9 53.8 55.4
URLCorpus None 47.0 42.4 39.7 47.8 73.2
Prompt 61.4 77.4 76.9 78.2 82.3
Reranking OLMoE Ask None 41.3 41.0 39.0 41.4 47.1
Prompt 43.4 51.9 49.9 50.2 51.6
SciDocs None 45.5 46.3 46.9 50.8 59.1
Prompt 53.6 69.6 73.1 75.1 77.2
StackOver None 25.8 24.8 20.1 26.4 33.2
Prompt 28.1 32.5 30.0 34.3 36.6
STS OLMoE Biosses None 39.3 13.6 7.7 29.7 64.0
Prompt 51.2 61.8 67.6 70.2 75.2
SickR None 50.3 46.3 26.4 53.0 58.9
Prompt 51.9 65.7 37.6 66.1 66.5
STS12 None 40.1 8.6 11.1 37.8 58.2
Prompt 51.3 53.8 37.5 63.6 66.4
STS13 None 40.5 21.1 18.2 43.4 61.5
Prompt 52.5 66.5 40.4 72.7 76.4
STS14 None 29.5 13.4 13.3 31.7 52.9
Prompt 41.1 56.8 33.9 64.2 68.2
STS15 None 30.8 27.8 22.5 33.3 63.2
Prompt 46.4 69.3 38.4 66.4 72.5
STS16 None 46.5 38.9 28.9 45.8 60.7
Prompt 52.4 70.1 49.4 68.3 71.8
STSBen None 42.2 23.4 17.5 44.5 57.4
Prompt 48.6 63.6 48.9 70.7 72.1
Summarization OLMoE Medrxiv None 28.4 20.9 21.1 29.8 31.8
Prompt 25.6 28.9 29.7 30.4 30.5
Table 7: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB tasks using the OLMoE model. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification Qwen1.5-MoE-A2.7B Emotion None 27.2 33.9 30.1 34.3 37.1
Prompt 37.0 48.5 47.4 47.2 51.3
Toxic None 53.0 61.1 21.3 52.9 61.4
Prompt 53.4 64.5 19.8 54.1 65.5
Tweet None 51.1 55.9 25.1 55.9 58.2
Prompt 56.1 61.1 38.6 60.7 62.3
Clustering Qwen1.5-MoE-A2.7B Medrxiv None 15.3 23.3 21.3 23.0 25.3
Prompt 14.2 24.6 19.8 21.8 27.9
20Groups None 12.0 31.5 25.1 27.4 34.4
Prompt 14.4 43.8 38.6 38.4 47.0
Pair Classification Qwen1.5-MoE-A2.7B SemEval None 42.0 38.8 34.7 42.5 47.5
Prompt 47.0 52.4 46.3 52.4 57.7
URLCorpus None 49.8 54.9 52.1 60.6 64.8
Prompt 56.7 68.7 65.8 68.2 75.5
Reranking Qwen1.5-MoE-A2.7B Ask None 43.1 45.8 43.4 47.3 49.1
Prompt 43.3 48.3 49.1 49.5 52.8
SciDocs None 49.6 60.6 55.3 67.0 71.6
Prompt 50.9 60.1 55.8 68.7 73.0
StackOver None 26.2 29.5 26.2 31.1 38.5
Prompt 28.8 31.3 30.2 35.2 44.5
STS Qwen1.5-MoE-A2.7B Biosses None 33.8 32.5 34.7 49.6 66.6
Prompt 55.1 55.8 48.5 68.4 74.5
SickR None 51.0 55.5 40.4 61.0 63.5
Prompt 50.2 59.7 51.1 64.3 69.1
STS12 None 40.2 16.9 18.6 46.3 52.3
Prompt 49.3 25.0 31.8 59.2 62.5
STS13 None 38.1 42.9 44.2 56.7 67.3
Prompt 53.3 57.5 54.6 73.4 75.6
STS14 None 28.1 26.5 25.6 45.4 53.7
Prompt 40.4 38.8 40.7 60.0 64.8
STS15 None 34.8 40.5 38.4 46.1 56.1
Prompt 40.7 52.3 54.2 58.8 66.8
STS16 None 47.6 51.0 48.1 58.1 64.4
Prompt 51.6 64.2 65.1 65.7 68.7
STSBen None 37.0 37.7 34.7 50.9 61.6
Prompt 45.6 47.8 54.5 64.5 70.0
Summarization Qwen1.5-MoE-A2.7B Medrxiv None 28.3 13.4 15.1 31.2 40.0
Prompt 27.0 23.0 21.9 27.3 31.0
Table 8: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with Qwen1.5-MoE-A2.7B models. The best result for each row is highlighted in bold.
Figure 7: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across across MTEB Tasks with PromptEOL [28]. The best result for each row is highlighted in bold.
Category Model Dataset Setting Router TC EC MoEE USMoE
Classification DeepSeekMoE-16B Emotion None 26.1 27.4 26.5 27.6 31.3
Prompt 37.9 48.3 46.1 46.4 52.6
Toxic None 53.3 60.4 58.1 53.1 61.7
Prompt 53.1 62.4 62.5 53.6 67.5
Tweet None 51.0 51.9 49.5 52.6 55.2
Prompt 54.9 58.4 57.5 58.9 61.0
Clustering DeepSeekMoE-16B Medrxiv None 15.1 23.0 17.3 22.0 25.7
Prompt 17.0 25.7 20.9 24.0 27.9
20Groups None 11.7 13.2 9.7 13.7 16.2
Prompt 18.6 32.3 19.8 33.0 37.6
Pair Classification DeepSeekMoE-16B SemEval None 44.6 40.2 32.6 43.5 44.6
Prompt 48.4 47.2 46.6 51.3 55.7
URLCorpus None 46.4 41.7 41.6 48.6 62.4
Prompt 66.5 72.4 61.1 75.4 80.0
Reranking DeepSeekMoE-16B Ask None 41.7 41.1 40.1 42.3 44.9
Prompt 43.5 43.8 44.7 46.9 50.6
SciDocs None 48.2 50.6 44.7 57.1 61.9
Prompt 58.3 65.6 55.3 72.6 72.9
StackOver None 25.7 24.9 20.4 27.3 30.2
Prompt 29.7 27.6 22.6 32.3 33.6
STS DeepSeekMoE-16B Biosses None 29.5 31.7 27.7 26.8 56.6
Prompt 47.0 40.1 41.5 57.6 67.2
SickR None 50.4 47.4 29.4 53.1 60.1
Prompt 56.0 61.9 38.7 65.8 68.0
STS12 None 44.0 4.3 13.9 45.0 48.6
Prompt 57.8 31.0 28.4 64.0 65.9
STS13 None 36.0 28.4 27.5 41.1 50.7
Prompt 55.3 56.0 41.2 70.9 76.1
STS14 None 25.4 12.0 13.0 28.2 41.2
Prompt 44.9 41.0 31.1 58.6 65.1
STS15 None 34.8 33.9 25.6 38.7 45.0
Prompt 49.7 46.5 33.0 58.5 64.1
STS16 None 44.9 34.4 33.1 46.9 55.2
Prompt 56.7 58.0 44.2 64.5 67.4
STSBen None 36.6 18.3 15.8 42.1 51.9
Prompt 54.9 57.7 39.0 67.8 71.0
Summarization DeepSeekMoE-16B Medrxiv None 24.9 22.0 18.5 24.4 29.9
Prompt 29.1 24.4 25.7 29.2 30.7
Table 9: Performance comparison of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks with DeepSeekMoE-16B models. The best result for each row is highlighted in bold.
(a) Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on Classification Task.
(b) Results distribution of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB Tasks
Figure 8: Illustration of comparing the performance of USMoE, Token Choice (TC), Expert Choice (EC), and MoEE across MTEB tasks and three SMoE models. Each benchmark is run 10 times, reporting both the mean and standard deviation to highlight the performance and stability of our method compared to the baselines.
Model Dataset TC EC USMoE-Sequence USMoE-Batch
DeepSeekMoE-16B Emotion 27.4 26.5 27.8 27.4
Toxic 60.4 58.1 60.1 59.2
Tweet 51.9 49.5 52.5 51.7
Table 10: Comparison of USMoE, Token Choice (TC), and Expert Choice (EC) on the classification task. Higher values are better, with the best results highlighted in bold.
Model Dataset $\alpha$
0.0 0.3 0.5 0.7 0.9 1.0
DeepSeekMoE-16B Emotion 27.4 27.1 27.8 27.6 27.7 26.5
Toxic 60.4 60.0 60.1 56.8 57.3 58.1
Tweet 51.9 53.2 52.5 53.3 52.9 49.5
Table 11: Performance comparison of DeepSeekMoE-16B across different classification datasets with varying $\alpha$ values. Higher is better; best results are in bold.
A.4.3 Training from scratch
Large models training.  USMoE not only delivers strong performance in base model training but also remains highly competitive at a large scale. Table 12 presents perplexity (PPL) results on the WikiText-103 and One Billion Words datasets using a large Transformer-XL model with 15 SMoE layers, 16 experts, and 420M parameters. The performance gap between USMoE and the baselines becomes even more pronounced at this scale, highlighting its strong scalability with increasing model complexity. Regardless of backbone size or the number of activated experts, USMoE consistently outperforms all baselines, demonstrating its effectiveness in scaling up large language models.
Transformer-XL(420M) WikiText-103 lm1b
Topkk TC EC USMoE TC EC USMoE
1 31.70 35.52 25.48 58.65 65.43 56.90
2 22.42 23.30 22.06 44.56 43.39 40.53
4 23.57 23.60 22.65 45.52 43.70 40.90
8 24.20 24.37 22.88 46.36 44.22 43.24
Table 12: Large Scale performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with perplexity on the WikiText-103 and One Billion Word test sets. Lower values are better, with the best results highlighted in bold.
[…]
USMoE (Topkk=2) 7.7620 81.5 40.1 88.5 87.8
       (Topkk=1.5) 6.6753 83.8 39.6 88.3 83.0
TC (Topkk=2) SMoE 7.7620 77.1 35.1 84.4 69.2
SMoE-DR 78.6 34.4 83.5 66.7
XMoE 76.7 35.3 83.3 67.4
StableMoE 77.7 34.3 83.9 60.8
EC (Topkk=2) 7.7620 81.5 39.3 88.0 75.6
Table 13: Accuracy performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) after fine-tuned on various datasets. Higher is better, best results are in bold.
Training from scratch.  To assess the effectiveness of our method, we compare USMoE with the Token Choice approaches, including SMoE [27], SMoE-Dropout (abbreviated as "SMoE-DR"), XMoE [10], and StableMoE [13], as well as the Expert Choice approach [58] for pre-training and fine-tuning tasks. We follow the approach of [7] and use a base Transformer-XL [14] with four decoder layers. We train both base and large-scale versions of Transformer-XL on four datasets (Enwik8, Text8, Wikitext-103, and One Billion Words) for 100k iterations, following the implementation in [7]. Then we fine-tune the pre-trained weights for text classification tasks, including SST-2 [49], SST-5 [49], IMDB [36], and BANKING77 [4]. More implementation details and additional results are provided in the Appendix A.
Fine-tuning.  We report the results of the fine-tuning experiment on the SST-2, SST-5, IMDB, and BANKING77 datasets in Table  13, using Transformer-XL pre-trained on enwik8. Overall, USMoE consistently achieves higher accuracy compared to other baselines across all datasets. The results demonstrate that our method is not only effective for pre-training tasks but also performs effectively on existing pre-trained models.
A.5 In-depth Analysis
We visualize the router behavior of USMoE in Figure 11 and contrast it with the router behaviors of the Token Choice approach (Figure 9) and the Expert Choice approach (Figure 10). Notably, the router in the OLMoE-1B-7B model exhibits a strong preference for specific experts. For instance, in the Emotion Classification task, Experts 8, 30, and 58 are consistently prioritized in both the Token Choice and Expert Choice approaches. This bias limits the model’s adaptability and effectiveness for downstream tasks. USMoE tackles this challenge by introducing the Unified Mechanism, which promotes more balanced and diverse expert selections, as illustrated in Figure 11. This enhancement enables USMoE to outperform the baselines on the Emotion Classification task.
Figure 9: Token Choice Router visualization for the OLMoE-1B-7B model on the Emotion Classification task. The scores of selected experts are replaced with -10.0 (lower than the minimum score) to enhance visualization. Best viewed in color.
Figure 10: Expert Choice Router visualization for the OLMoE-1B-7B model on the Emotion Classification task. The scores of selected experts are replaced with -10.0 (lower than the minimum score) to enhance visualization. Best viewed in color.
Figure 11: USMoE Router visualization for the OLMoE-1B-7B model on the Emotion Classification task. The scores of selected experts are replaced with -10.0 (lower than the minimum score) to enhance visualization. Best viewed in color.
We track the number of unique experts utilized by the OLMoE-1B-7B model for each sequence in the Emotion Classification task as Figure 12(a). Our analysis reveals that the Expert Choice approach employs 11 out of 16 experts, indicating a lower level of specialization among experts. In contrast, both USMoE and the Token Choice approach use an average of 0.9 to 1 expert per sequence, demonstrating superior expert specialization. Furthermore, we analyze the token dropping behavior of the Expert Choice approach and observe a significant increase in dropping rates when scaling to larger datasets or models, such as pre-training the Transformer-XL Large model on the One Billion Word dataset, as shown in Figure 12(b). This increase in dropping rates may negatively impact model performance. In contrast, our method maintains a consistently low dropping rate (<0.1), demonstrating its superiority over the Expert Choice approach for scalability. Additionally, our method proves more robust than the Token Choice approach, as it effectively drops irrelevant tokens without compromising performance.
(a) Number Experts per Sequence of USMoE, TC, and EC on Emotion dataset.
(b) Token Dropping of USMoE, Token Choice (TC), Expert Choice (EC) for Pre-training on One Billion Word dataset.
Figure 12: Comparison of the number of experts per sequence for USMoE, Token Choice (TC), and Expert Choice (EC) on the Emotion dataset using the OLMoE-1B-7B model, along with a comparison of token dropping rates for USMoE, TC, and EC during pre-training on the One Billion Word dataset.
A.6 Implementation Details
For the Without Training experiments, we implement our method based on the publicly available MoEE implementation [34] [^1] https://github.com/tianyi-lab/MoE-Embedding. Due to resource constraints, we validate our method and the baselines using 4-bit quantization with a batch size of 128. For the OLMoE-1B-7B model, we conduct experiments on a single H100 GPU, while for the Qwen1.5-MoE-A2.7B and DeepSeekMoE-16B models, we utilize two H100 GPUs.
The base Transformer-XL variant [7] comprises four Transformer decoder layers, each with an input dimension of 256. Each layer includes a self-attention mechanism with eight attention heads, followed by a Feed-forward Neural Network (FFN) that has an inner dimension of 512. The dropout ratio is set at 0.1. We divide the FFN into 16 experts, each with the same dimensions. For the larger variants, we scale the model up to twelve layers.
Our experiments are based on the publicly available SMoE-Dropout implementation [7] [^2] https://github.com/VITA-Group/Random-MoE-as-Dropout. The pre-training experiments were conducted using a single H100 GPU, while the fine-tuning experiments were performed on a single A100 GPU. It is important to note that parallel training on multiple GPUs may produce different results.
A.6.1 Pre-training Experiments
Transformer-XL(20M) Enwik8 Text8 WikiText-103 lm1b
USMoE (Topkk=2) 1.18 1.20 29.20 56.90
       (Topkk=1.5) 1.19 1.28 30.67 57.55
TC (Topkk=2) SMoE 1.20 1.29 30.16 58.00
SMoE-DR 1.56 1.56 58.37 93.17
XMoE 1.21 1.28 30.34 58.33
StableMoE 1.20 1.28 29.97 58.25
EC (Topkk=2) 1.18 1.24 29.83 58.60
Table 14: Performance comparison of USMoE, Token Choice (TC), and Expert Choice (EC) across multiple datasets, with BPC on the Enwik8 and Text8 test sets, and perplexity on the WikiText-103 and One Billion Word test sets. Lower values are better, with the best results highlighted in bold.
We provide the USMoE implementation details for pre-training our Transformer-XL base and large on enwik8, text8, WikiText-103, and One Billion Word in Table 15.
Table 15: Implementation details for pre-training experiments on enwik8, text8, WikiText-103, and One Billion Word datasets.
Dataset Input length Batch size Optimizer Lr # Iterations
enwik8 512 48 Adam 2.5e-4 100k
text8 512 48 Adam 2.5e-4 100k
WikiText-103 512 22 Adam 2.5e-4 100k
One Billion Word 512 11 Adam 2.5e-4 100k
A.6.2 Fine-tuning Experiments
To perform the fine-tuning experiments, we utilize the same model architecture as in the pre-training phase. Table 16 presents the implementation details for the fine-tuning experiments conducted across four different datasets.
Table 16: Implementation for fine-tuning experiments on downstream tasks.
Dataset Input length Batch size Optimizer Lr # Epochs
SST-2 512 16 Adam 1e-4 15
SST-5 512 16 Adam 1e-4 15
IMDB 512 4 Adam 1e-4 15
BANKING77 512 16 Adam 1e-4 15
Limitations
Our study focuses on enhancing the efficiency and effectiveness of Large Language Models (LLMs) through SMoE. Our approach proves effective in both training and non-training settings. While the results are promising, our pre-training experiments were constrained by computational resources, limiting us to medium-scale datasets and a base Transformer-XL model. Consequently, further empirical evaluation is required to assess the scalability of USMoE beyond 100B parameters and compare it with other SMoE strategies in modern LLMs.
Ethics Statement
Despite encouraging results, training large-scale LLMs remains highly resource-intensive, requiring careful management of computational costs. Additionally, our study relies on web-sourced data, which is known to contain gender and racial biases, highlighting the need for further efforts to mitigate these issues. Lastly, while our work represents a significant step toward advancing LLMs development, it also emphasizes the importance of robust regularization to prevent potential misuse in harmful applications.